OpenStudy (anonymous):

Solve. 3-0.5(u-9)=0.6(10u-9)

OpenStudy (anonymous):

\[3-0.5u+4.5 = 6u-5.4\] \[-1.5u+7.5=6u-8.4\] \[-6.5u = -12.9\] \[u = 1.988~\]